The data below represent commute times (in minutes) and scores on a well-being survey. Commute Time (minutes), x 5 20 25 40 50 84 105 Well-Being Index Score, y 69.2 68.0 67.4 66.6 66.2 65.1 63.3 Given that r = -0.9841, Sx = 35.981, Sy = 1.942.

a) Find the equation for the regression line.

b) Interpret the slope and y-intercept.

c) Find the predicted value of index score when x = 40. Find the residual for the value of y =66.6 and interpret it.
Answer #1

a)

S.No х 1 2 3 4 5 6 7 Total 5 20 25 40 50 84 105 329 69.2 68 67.4 66.6 66.2 65.1 63.3 465.8 1764.0000 729.0000 484.0000 49.000

Least square line equation: ŷ =69.04-0.0531*x

b)

slope -0.0531 indicate that for 1 unit increase in commute time , Well-Being Index Score decreases by 0.0531 on average

intercept 69.04 indicate that for 0 minute commute time ; a person predicted Well-Being Index Score is 69.04

c)

predicted val=69.04+40*-0.0531= 66.92

residual =actual-predicted value =66.6-66.92 = -0.32

this tells us that actual value is 0.32 less than what we predict from the regression line,